for the bottom two the equation that i used for the first two didn'tExercise 1-8 Product Costs and Period Costs; Variable and Fixed Costs [LO1-3, LO1-4] Kubin Companys relevant range of produc work. what numbers are you supposed to use to get this. I dont even know where to start
Answer #1
PRODUCT COST:
Product Cost includes :
Direct Materials
Direct Labor
Manufacturing Overhead
Note: Fixed cost remains constant
A B=A*17250 C=A*20500
Average Variable Costs for
Cost per unit 17250 units 20500 units
x Direct Materials $7.50 $129,375 $153,750
y Direct Labor $4.50 $77,625 $92,250
x Variable Manufacturing Overhead $2.00 $34,500 $41,000
D=x+y+z Total Variable Production costs $241,500 $287,000
E Fixed Manufacturing Overhead($5.50*17250) $94,875 $94,875
F=D+E Total amount of product cost $336,375 $381,875
Total amount of product cost to make 17250 units $336,375
Total amount of product cost to make 20500 units $381,875
PERIOD COST:
Period Cost includes:
Selling expense
Administrative expense
Sales Commission
A B=A*17250 C=A*20500
Average Variable Costs for
Cost per unit 17250 units 20500 units
x Sales Commission $1.50 $25,875 $30,750
y Variable Administrative expense $1.00 $17,250 $20,500
D=x+y Total Variable Period costs $43,125 $51,250
E Fixed Selling Expense($4*17250) $69,000 $69,000
F Fixed Administrative Expense($3*17250) $51,750 $51,750
G=D+E+F Total amount of Period cost $163,875 $172,000
Total amount of product cost to make 17250 units $163,875
Total amount of product cost to make 20500 units $172,000

Total Period Cost of 14000 units

Sales Commission =1.5*14000=$21,000

Variable Administrative Expense =1*14000=$14,000

Fixed Selling Expense=$69,000

Fixed administrative Expense=$51,750

Total Period Cost =21000+14000+69000+51750=

$155,750